OpenStudy (anonymous):

show me how to graph the line for (4.3)(-4-5) and the given pair of points to find slope

OpenStudy (lgbasallote):

slope = y2 - y1/ x2 - x1 m = -5 - 3/ -4 -4 m = -8/0

OpenStudy (phi):

I assume the 2 points are (4,3) and (-4,-5)  To find the equation between 2 points, first find the slope (called m) the slope is the distance in y divided by the distance in x \[m=\frac{y_2-y_1}{x_2-x_1}\] Here we get \[m=\frac{-5 - 3}{-4-4}= \frac{-8}{-8}= 1\] Now use m in the formula \[y-y_1= m(x-x_1)\] where x1,y1 is one of the points (either one will do). Let's use (4,3): y-3= 1*(x-4) we get y= x-4+3 or y= x-1